OpenStudy (unklerhaukus):

\[\sqrt{a+ib}\cdot\sqrt{\frac1{a+ib}}=\begin{cases} ~\\~\\~\end{cases}\]

mathslover (mathslover):

first step should be square both sides

OpenStudy (unklerhaukus):

but what is the RHS? @mathslover

mathslover (mathslover):

let RHS be x

OpenStudy (unklerhaukus):

how do you know there is only one solution?

mathslover (mathslover):

oh k wait i have an easier way

OpenStudy (anonymous):

it is one from your eyeballs

OpenStudy (unklerhaukus):

@satellite73 how do you know there is only one solution?

OpenStudy (anonymous):

not sure what you mean

OpenStudy (unklerhaukus):

well my eyes see that if a=b=0 then the is no solution

mathslover (mathslover):

\[\large{\frac{(a+ib)(a-ib)}{a^2+b^2}=x^2}\] \[\large{\frac{a^2+b^2}{a^2+b^2}=x^2}\] \[\large{1=x^2}\] \[\large{x=1}\]

OpenStudy (anonymous):

\[\sqrt{z}\times\frac{1}{\sqrt{z}}=\sqrt{\frac{z}{z}}=\sqrt{1}=1\]

mathslover (mathslover):

that is your answer

mathslover (mathslover):

right satellite :D

OpenStudy (anonymous):

\[\sqrt{a+ib}\cdot\sqrt{\frac1{a+ib}}=\begin{cases} ~\\~\\~\end{cases}\] Which is equal to \[\sqrt{\frac{ a+ib }{ a+ib }} = \sqrt{1} = \pm 1\]

mathslover (mathslover):

\pm matters a lot

mathslover (mathslover):

got it @UnkleRhaukus ?

mathslover (mathslover):

x = \(\pm 1\)

OpenStudy (anonymous):

A lot of lot.

mathslover (mathslover):

unklerhaukus is typing a reply :)

OpenStudy (unklerhaukus):

\[\sqrt{a+ib}\cdot\sqrt{\frac1{a+ib}}=\begin{cases} 1&...\\0& a=b=0\\-1&...\end{cases}\]

OpenStudy (anonymous):

depends on whether you think \(\sqrt{a}\) means the the positive root of \(a\) square root is a well defined function, or whether you think \(\sqrt{a}\) means any solution to \(x^2=a\) i am of the former view, but that is just my opinion

OpenStudy (anonymous):

really? what number is \(\frac{1}{0}\)?

OpenStudy (unklerhaukus):

whops,

OpenStudy (unklerhaukus):

\[\sqrt{a+ib}\cdot\sqrt{\frac1{a+ib}}=\begin{cases} 1&...\\ ? & a=b=0\\-1&...\end{cases}\]

OpenStudy (anonymous):

It is still 0 if a=b=0.

mathslover (mathslover):

yes but that is not possible

OpenStudy (anonymous):

In fact no, it's still \pm 1 if a=b=0

OpenStudy (anonymous):

\(\frac{0}{0}=1\) ?

mathslover (mathslover):

I think the given problem needs something more info. : \[\large{\textbf{IF} a \space \textbf{and} \space b\ne 0}\]

OpenStudy (anonymous):

No... they cancel.

OpenStudy (anonymous):

\[\frac{0}{0}=\frac{\cancel{0}}{\cancel{0}}=1\]

OpenStudy (anonymous):

\[a \times \frac{ b }{ a }\] = b. We can all agree on that can't we? Just because a can take the value zero doesn't mean that there is a different solution there. The fact remains that the a's cancel.

OpenStudy (anonymous):

i like cancelling zeros

mathslover (mathslover):

0/0 = undefined

OpenStudy (anonymous):

but even in the world of complex numbers, you cannot divide by 0

OpenStudy (anonymous):

But you aren't dividing by 0, it has been cancelled.

OpenStudy (anonymous):

implicit is writing \(\frac{1}{z}\) is that \(z\neq 0\)

OpenStudy (anonymous):

Yes this should have been written in the question really.

OpenStudy (anonymous):

no it is in there already

mathslover (mathslover):

guys let unkle rhaukus think also :D

OpenStudy (unklerhaukus):

ok from now on Z≠0

OpenStudy (anonymous):

i guess i ldepends on how you view this symbol \[\sqrt{a}\]

OpenStudy (unklerhaukus):

yeah i have been thinking about it today http://openstudy.com/users/unklerhaukus#/updates/502f3359e4b0ac2883165a00

OpenStudy (anonymous):

to some, myself included, a real number has two "square roots" the two square roots of \(a\) are \(\sqrt{a}\) and \(-\sqrt{a}\) meaning the symbol \(\sqrt{a}\) means the "positive square root" i.e. the positive numbers whose square is \(a\)

OpenStudy (unklerhaukus):

\[\sqrt{a+ib}\cdot\sqrt{\frac{1}{a+ib}}\]\[=\sqrt{a+ib}\cdot\sqrt{\frac{1}{a+ib}\times\frac{a-ib}{a-ib}}\]\[=\sqrt{a+ib}\cdot\sqrt{\frac{a-ib}{a^2+b^2}}\]

OpenStudy (anonymous):

we can try a little experiment and see what the wolf gods think http://www.wolframalpha.com/input/?i=sqrt%282%2B3i%29*1%2F%28sqrt%282%2B3i%29%29

OpenStudy (unklerhaukus):

dont you mean http://www.wolframalpha.com/input/?i=sqrt%282%2B3i%29*sqrt%281%2F%282%2B3i%29%29, different result

mathslover (mathslover):

that is what i did @UnkleRhaukus

OpenStudy (unklerhaukus):

Computation timed out. @satellite73

OpenStudy (anonymous):

mine gave me 1 instantly

OpenStudy (anonymous):

i used your link and got one as well did it time out for you?

OpenStudy (unklerhaukus):

well it does say one 1, it also say Computation timed out. which explains the missing solutions , I have been trying to prove (11) http://mathworld.wolfram.com/SquareRoot.html

OpenStudy (anonymous):

it is graphing it for you in two dimensions, as if you are in the \((a,b)\) plane but you notice that it is 1 straight across

OpenStudy (unklerhaukus):

yeah i see that bit of the graph but how did can i arrive at that

OpenStudy (anonymous):

ok look all the way at the bottom you will get a better answer than i provided

OpenStudy (unklerhaukus):

the scary looking ∑ sums?

OpenStudy (anonymous):

interesting, wolfram is contradicting itself here check this out http://www.wolframalpha.com/input/?i=sqrt%28-2-3i%29*sqrt%281%2F%28-2-3i%29%29%2C

OpenStudy (unklerhaukus):

hmm machine problem , , how do they get the argument thing \(\text{Arg(a+ib)}\)

OpenStudy (anonymous):

\[ = \left\{\begin{array}{rcc} -1 & \text{if} & arg(a+bi)\geq \pi \\ 1& \text{othewise} \end{array} \right. \]

OpenStudy (unklerhaukus):

is that from a picture ?